# Identifying Aspects in Peer Reviews

### v1

@@ Abstract
Peer review is central to academic publishing, but the growing volume of submissions is straining the process. This motivates the development of computational approaches to support peer review. While each review is tailored to a specific paper, reviewers often make assessments according to certain aspects such as Novelty, which reflect the values of the research community. This alignment creates opportunities for standardizing the reviewing process, improving quality control, and enabling computational support. While prior work has demonstrated the potential of aspect analysis for peer review assistance, the notion of aspect remains poorly formalized. Existing approaches often derive aspect sets from review forms and guidelines of major NLP venues, yet data-driven methods for aspect identification are largely underexplored. To address this gap, our work takes a bottom-up approach: we propose an operational definition of aspect and develop a data-driven schema for deriving fine-grained aspects from a corpus of peer reviews. We introduce a dataset of peer reviews augmented with aspects and show how it can be used for community-level review analysis. We further show how the choice of aspects can impact downstream applications, such as LLM-generated review detection. Our results lay a foundation for a principled and data-driven investigation of review aspects, and pave the path for new applications of NLP to support peer review. [^1] Our code and data are available at https://anonymous.4open.science/r/aspects-in-reviews-737E/.
@@ 1 Introduction
Peer reviews are the central component of the reviewing process. While individual review texts can vary widely, reviewers within the same research community tend to focus on a specific set of general quality categories, or aspects, such as Methodology and Novelty. These aspects can be found in review forms, instructional materials, guidelines, and the resulting review texts. Aspects allow fine-grained comparison of submissions across different dimensions of quality. A comprehensive set of aspects shared among reviewers is critical for ensuring review quality and consistency, and prior work has demonstrated the potential of aspect-based tools to support the review writing process Sun et al. (2024a, b).
Figure 1: An aspect is a characteristic of a paper that a reviewer makes a judgment on. In this paper, we leverage an LLM to derive aspects from a corpus of peer reviews.
@@
Yet, several open questions remain. First, what is an aspect? Most prior work derives aspects from review forms and guidelines. The lack of an operational definition of aspect prevents the comparison of aspect schemata across studies. Second, what is the optimal level of aspect granularity? While prior work operates with top-down, coarse-grained aspect schemata derived from review forms and guidelines, it is unclear whether these schemata are comprehensive or provide sufficient granularity for NLP applications. Third, how can fine-grained aspect analysis support peer review? While aspects have been applied to certain tasks, the tasks that require a higher level of granularity are underexplored.
@@
To address these questions, this work introduces an alternative, data-driven approach to peer review aspect analysis. We propose an operational definition of aspect grounded in its role within the evaluation process. We develop a semi-automatic approach that leverages a state-of-the-art large language model (LLM) to identify aspects in reviews, and apply it to a large collection of peer reviews to extract aspects in a bottom-up fashion (see Figure 1). Building on these results, we develop a multi-level taxonomy of aspects and present a novel dataset of peer reviews augmented with their corresponding aspects. Our dataset facilitates the exploration of two tasks: predicting the review aspects that should be focused on given a paper (paper aspect prediction), and identifying the aspects that a review focuses on (review aspect prediction). Based on these tasks, we conduct a detailed empirical study of aspect at the community level. Furthermore, we demonstrate that a comprehensive, fine-grained set of aspects allows for a new dimension in comparing reviews, offers a nuanced assessment of review quality in terms of specificity, and helps the detection of automatically generated reviews. Our work paves the path for data-driven, bottom-up analysis of aspects in peer reviews and enables new applications of NLP to support review writing. To summarize, this work contributes the following:
@@
  • We propose an semi-automatic approach which leverages a state-of-the-art LLM to derive a comprehensive set of aspects from peer reviews in a bottom-up fashion.
@@
  • We evaluate models on two tasks: predicting review aspects that should be focused on given a paper, and identifying the aspects covered in a review.
  • We show how aspect analysis provides new insights into the reviewing process and how a finer-grained set of aspects helps in detecting LLM-generated reviews.
@@ 2.2 Aspect in peer review
Beyond their role in review forms and guidelines, aspects are utilized in several other contexts related to peer review. Aspects are used in analyzing the sentiment of reviews and discovering the discourse relations within reviews Chakraborty et al. (2020); Kennard et al. (2022). In automatic review generation, Wang et al. (2020) generate reviews using aspect-based knowledge graphs, and Gao et al. (2024) prompt LLMs with aspect-based questions to generate reviews. Aspects are included in supporting systems for review writing, which have been shown to improve the comprehensiveness of the reviews written by reviewers with varying levels of experience Sun et al. (2024a, b). Table 1 summarizes the aspects used in the studies cited in this section. In some other review guidelines, such as the ARR review form and ACL’23 review policies, aspects are presented in an “informal” manner – they are listed as examples rather than parts of a comprehensive checklist. As noted by Kuznetsov et al. (2024), review guidelines for major NLP venues often rely on coarse-grained aspects and they lack comprehensiveness. Our work advances the study of aspects in peer review by exploring an alternative, data-driven approach to deriving finer-grained and more comprehensive aspect sets.
@@
Clarity, Impact/Motivation, Meaningful Comparison, Originality, Replicability, Soundness/Correctness, Substance, Summary
@@ 3.1 Definition of aspect
While aspects are outlined in review guidelines and used in related work, a formal definition of aspect is lacking. To address this, we propose an operational definition of aspect as a characteristic of a paper that a reviewer makes a judgment on when evaluating the paper, which is later used to compare the manuscripts to each other and to the quality standards provided by the review guidelines. From this definition, we do not consider terms such as Acceptance Decision and Summary as aspects, since they are not characteristics of a paper. An aspect can be general, such as Soundness, or specific, such as Missing Citations on Controlled Generation. However, aspects that are either too general or too specific are difficult to use for comparing manuscripts to each other or to the publication standards. Therefore, an aspect taxonomy that accommodates different levels of granularity is needed.
@@ 3.2 Method
To establish a comprehensive set of aspects, we selected paper reviews from NLP and machine learning (ML) conferences across different time periods. We randomly selected 50 papers from each of the NLPeer Dycke et al. (2023) and EMNLP23222Publicly available through the OpenReview API. datasets. We used the keywords “natural” and “language” to filter NLP-related papers from ICLR. We randomly selected 50 NLP-related papers from each of the ICLR conferences from 2020 to 2024. We selected 350 papers in total, corresponding to 1094 reviews. We segmented the reviews into sentences using NLTK Punkt Tokenizer Bird and Loper (2004) and performed the identification at the sentence level. See Appendix B for more details regarding the data.
@@
We used OpenAI GPT-4o [^3] Version 2024-08-06. Results were obtained between Oct 31 and Nov 20, 2024. to identify aspects from the reviews. The prompt we used is shown in Table 7. Since the identification was performed in an unsupervised setting, the identified aspects appear inconsistent, with variations like Result and Results. We post-processed the results to reduce such variations. We first identified the most frequently occurring aspects in the results, which were used as keywords to categorize and match the remaining ones. We grouped related terms together, such as Clarification and Clarity. We removed terms that we considered to be too general, such as Weakness, Strength, Question, and Comment. For the terms that cannot be matched, we omitted those that appear less than 50 times in the annotations.
We notice that our post-processing method may lead to unrelated terms being grouped together. For example, using Improvement as a keyword groups Performance Improvement and Improvement Recommendation (i.e., suggestions to improve paper quality) together. We manually checked the categorization results to verify and correct any inappropriate groupings.
In cases where a single term contains multiple aspects, it is placed into more than one category. For example, Comparison with Related Work is included in both Comparison and Related Work categories.
@@ 3.3 Results
Some aspects are not present in our taxonomy as review forms for major NLP venues have dedicated fields for their evaluation, making them appear much less frequently in the review text (e.g., Reproducibility). See Table 8 for examples of the dataset we created which pairs aspects with real-life review sentences and Table 9 for the aspect frequencies. In the following sections of the paper, we use both the coarse and fine label sets in our experiments and analyses.
@@ 3.4 Validity check
We conducted a human evaluation to verify the model annotations, involving three human annotators to evaluate the LLM annotations mapped to the coarse label set. We asked the annotators to determine whether a review sentence addresses the aspects identified by GPT-4o (see Table 14 for examples of the questionnaire entries). We observe that on average human annotators agree with 91% of the LLM annotations, along with fair inter-annotator agreement as measured by Fleiss’ Kappa Fleiss (1971) (see Table 4). These results suggest that the model annotations largely align with human judgments, and that human annotators can effectively understand our taxonomy.
See Appendix B.2 for more details regarding the validity check.
@@ 4.1 Method
It is important to note that PAP is inherently more challenging than RAP. PAP operates as a heuristic method grounded in statistics, or in a more advanced setting (i.e., when the input it the full paper), goes beyond heuristics to infer strengths or weaknesses in certain aspects of the paper. In contrast, RAP resembles summarization, as it extracts aspects directly from the review text, making it comparatively less challenging than PAP.
We modeled both tasks as multi-label sequence classification. For both tasks, we tested bag-of-words with random forest (BoW+RF) and RoBERTa Liu et al. (2019). Both models are strong in supervised settings, and they are lighter-weight alternatives to LLMs. We used focal loss Lin et al. (2017) to address label imbalance. For PAP, we also experimented with GPT-4o in both zero-shot and few-shot settings. Our evaluation metrics include precision, recall, F1 score, and Jaccard similarity score. See Appendix C for more details regarding the experimental setup.
@@ 5.1 Aspect analysis
Table 5: The submission tracks with the 5 highest frequencies of Analysis in EMNLP23. Table 22 shows the full results.
@@ 5.2 Review comparison
To compare the reviews, we predicted the aspects in each review and calculated the Jaccard similarity between sets of aspects. Figure 3 visualizes the similarity between each pair of human-written and LLM-generated reviews generated using Liang et al. (2024b)’s prompt.
@@
Figure 3: The heatmap of the Jaccard similarity between each pair of the human-written reviews and LLM-generated reviews generated using EMNLP23 papers Liang et al. (2024b)’s prompt. Figure 7 in Appendix D.2 shows the rest of the results.
We obverve that LLM-generated reviews show a higher degree of similarity to each other in terms of aspects. This suggests that LLM-generated reviews may be more general than human-written ones, with the model tending to comment on similar sets of aspects across different papers. This could also imply that the quality of LLM-generated reviews is still lacking in terms of specificity.
@@ 5.3 LLM-generated review detection
Table 6 shows the detection results. We used Jaccard similarity and compared our method with an implementation of Equation 2 using Sentence-BERT (SBERT) Reimers and Gurevych (2019), which calculates the similarity between the embeddings of review texts. Experiments were conducted using both the coarse and fine label sets, as well as ACL’18 which is a commonly used aspect set in previous studies (see Table 1). We calculated a random baseline that selects a review at random as the LLM-generated review. Note that the fine aspect set consistently outperform ACL’18 and SBERT, and our method is robust across LLM-generated reviews generated using prompts with different levels of prompt engineering.
@@
Practically, the detection method introduced in this section does not perform as well as more common reference-based methods (e.g., Yu et al. (2024), which reports over 90% accuracy) or zero-shot detection using state-of-the-art LLMs (see GPT-4o results in Table 24), and the setting is specific, with only one LLM-generated review per paper. Nevertheless, our findings validate the aspects and the dataset curated in this work, and more importantly, they provide a key insight into LLM-generated reviews, that current LLMs tend to produce reviews that are general in terms of aspects, covering similar aspects across different papers.
@@ 6 Conclusion
In this paper, we introduce a data-driven approach to peer review aspect analysis. We provide an operational definition of aspect, and develop an semi-automatic approach using GPT-4o to identify aspects from peer reviews. We propose a taxonomy that involves a comprehensive set of aspects with different granularity, and introduce a new dataset of peer reviews augmented with aspects. We introduce two tasks, paper aspect prediction and review aspect prediction, and show how they contribute to a detailed empirical study of aspects. We further demonstrate that a comprehensive, fine-grained set of aspects opens a new dimension in review comparison and provides insights into the quality of automatically generated reviews.
@@
As discussed in Section 3.1, the concept of aspect is difficult to define. In this work, we adopted an operational definition as a practical approach. Defining individual aspects is also challenging. For instance, it is difficult to determine the scope of “Methodology.” Polysemy further complicates this issue–for example, “Improvement” may refer to either a method’s improvement, or places where the paper can be improved. We did a manual inspection of the categorization results to minimize the impact of these issues.
It is important to note that the categorization proposed in this work may not be optimal, and alternative approaches to categorizing aspects may exist. Our taxonomy is an attempt to streamline the analysis and application of aspects, and it has been shown to be effective for the purposes of this study.
@@
As reported in Section 3.4, GPT-4o shows some sensitivity to prompts and temperature, which may indicate uncertainty or inaccuracies in some annotations. Nevertheless, our human evaluation affirms the overall validity of the GPT-4o annotations. The applications of these aspect annotations, especially in the LLM-generated review detection task, provide further support for their reliability. Though we specify the seed (See Appendix B and C), the exact reproducibility of the results related to GPT-4o may still be a concern due to its closed source nature. We chose to use GPT-4o because it is among the best performing models across a wide range of tasks.
@@
For human annotation, we recognize that using aspect labels from our taxonomy may introduce potential biases. Asking the annotators to do an open-ended annotation task without pre-defined labels seems more straight forward, but it is very challenging in terms of the cognitive burden on annotators and annotation consistency. Automation bias is also a concern. We address this by asking annotators to provide explanations for their annotations.
All papers associated with the ICLR20-24 and EMNLP23 datasets are camera-ready versions. Though initial drafts are more preferable, we did not have access to them via OpenReview. This may affect the validity of the results of PAP experiments. However, we do not consider the use of camera-ready versions problematic in Section 5.2 and 5.3. Even if some issues raised by the reviewers were fixed in the camera-ready versions, the generated reviews are still expected to differ from each other. While the NLPeer dataset includes initial drafts, we did not use them for generating reviews due to the limited number of reviews per paper (around 2 reviews per paper).
@@
Another limitation of our analysis is the potential biases introduced by data imbalance regarding paper acceptance. The EMNLP23 dataset contains only 9 rejected papers, which leads to a very imbalanced distribution. Similarly, in the NLPeer dataset, 69% of the papers are accepted papers Dycke et al. (2023). Such skew may lead to an overestimation of the frequency of certain aspects. Since most papers in the EMNLP23 and NLPeer datasets are accepted papers, aspects commonly associated with accepted papers may appear more frequently than they would in a more balanced dataset. Consequently, some of our findings may be influenced by this bias. If these overrepresented aspects are used to inform review forms or guidelines, they may introduce new biases into the review process, which reduces their effectiveness. Despite this limitation, we chose these datasets because they are so far the most comprehensive review datasets in the domain of NLP.
@@
In Section 5.3, we only focused on end-to-end prompts as they represent common types of prompts for review generation. We did not consider prompts that involves paper-specific aspects, as we consider it to be human-in-the-loop prompts, where the user must first read the paper to identify paper-specific aspects and then incorporate them into the prompt. This approach represents a form of human-AI collaboration that may further assist the peer review process. We plan to explore this direction in future work.
@@
This work does not suggest or imply that the proposed task of predicting the aspects to focus on a given paper may replace human involvement. Instead, it is designed to assist reviewers by recommending aspects to consider. Reviewers retain full autonomy over their decisions regarding whether to include the suggested aspects.
Since all the review data used in this study are publicly available and anonymized, processing it with commercial LLMs such as OpenAI GPT-4o does not raise ethical concerns.

### v2

@@ Abstract
Peer review is central to academic publishing, but the growing volume of submissions is straining the process. This motivates the development of computational approaches to support peer review. While each review is tailored to a specific paper, reviewers often make assessments according to certain aspects such as Novelty, which reflect the values of the research community. This alignment creates opportunities for standardizing the reviewing process, improving quality control, and enabling computational support. While prior work has demonstrated the potential of aspect analysis for peer review assistance, the notion of aspect remains poorly formalized. Existing approaches often derive aspects from review forms and guidelines, yet data-driven methods for aspect identification are underexplored. To address this gap, our work takes a bottom-up approach: we propose an operational definition of aspect and develop a data-driven schema for deriving aspects from a corpus of peer reviews. We introduce a dataset of peer reviews augmented with aspects and show how it can be used for community-level review analysis. We further show how the choice of aspects can impact downstream applications, such as LLM-generated review detection. Our results lay a foundation for a principled and data-driven investigation of review aspects, and pave the path for new applications of NLP to support peer review. [^1] Our code and data are available at https://anonymous.4open.science/r/aspects-in-reviews-737E/.
@@ 1 Introduction
Peer reviews are the central component of the reviewing process. While individual reviews can vary widely, reviewers within the same community tend to focus on a specific set of general quality categories, or aspects, such as Clarity and Novelty. These aspects can be found in review forms, instructional materials, guidelines, and the resulting review texts. Aspects allow comparison of submissions across different dimensions of quality. A comprehensive set of aspects shared among reviewers is critical for ensuring review quality and consistency, and prior work has demonstrated the potential of aspect-based tools to support the review writing process Sun et al. (2024a, b).
@@
Yet, several open questions remain. First, what is an aspect? Most prior work derives aspects from review forms and guidelines. The lack of an operational definition of aspect prevents the comparison of aspect schemata across studies. Second, what aspect granularity is appropriate for different tasks? While prior work operates with top-down, coarse-grained aspect schemata derived from review forms and guidelines, it is unclear whether they are comprehensive or provide sufficient granularity for NLP applications. Third, how can fine-grained aspect analysis support peer review? While aspects have been applied to certain tasks, the tasks that require a higher level of granularity are underexplored.
@@
To address these questions, this work introduces an alternative, data-driven approach to peer review aspect analysis. We propose an operational definition of aspect grounded in its role within the evaluation process. We develop a semi-automatic approach that leverages a state-of-the-art large language model (LLM) to identify aspects in reviews, and apply it to a large collection of peer reviews to extract aspects in a bottom-up fashion. Building on these results, we develop a multi-level taxonomy of aspects and present a novel dataset of peer reviews augmented with their corresponding aspects. Our dataset facilitates the exploration of two tasks: predicting the review aspects that should be focused on given a paper (paper aspect prediction), and identifying the aspects that a review focuses on (review aspect prediction). Based on these tasks, we conduct a detailed empirical study of aspect at the community level. Furthermore, we demonstrate that a comprehensive, fine-grained set of aspects allows for a new dimension in comparing reviews, offers a nuanced assessment of review quality in terms of specificity, and can be used for detection of automatically generated reviews.
Our work paves the path for data-driven analysis of aspects and enables new NLP applications to support peer review. Our method offers a bottom-up perspective that complements existing top-down schemata, and can be integrated with them by, for example, allowing domain experts to refine LLM-derived aspects into higher-level, domain-specific categories. Together, these perspectives enable a more comprehensive aspect-based analysis in peer review. To summarize, this work contributes the following:
@@
  • We propose an semi-automatic approach to derive a comprehensive set of aspects from peer reviews in a bottom-up fashion.
@@
  • We evaluate models on two tasks: predicting aspects to be focused on given a paper, and identifying the aspects covered in a review.
  • We show that finer-grained aspect analysis offers new insights into and support for the peer review process.
@@ 2.2 Aspect in peer review
Beyond their role in review forms and guidelines, aspects are utilized in several other contexts related to peer review. Aspects are used in analyzing the sentiment of reviews and discovering the discourse relations within reviews Chakraborty et al. (2020); Kennard et al. (2022). In automatic review generation, Wang et al. (2020) generate reviews using aspect-based knowledge graphs, and Gao et al. (2024) prompt LLMs with aspect-based questions to generate reviews. Aspects are included in supporting systems for review writing, which have been shown to improve the comprehensiveness of the reviews written by reviewers with varying levels of experience Sun et al. (2024a, b). Table 1 summarizes the aspects used in these studies.
@@
Clarity, Impact/Motivation, Meaningful Comparison, Originality, Replicability, Soundness/Correctness, Substance
@@
In some other review guidelines, such as ARR and ACL’23, aspects are listed as examples rather than parts of a comprehensive checklist. As noted by Kuznetsov et al. (2024), review guidelines for major NLP venues often rely on coarse-grained aspects and they lack comprehensiveness. Our work advances the study of aspects by exploring an alternative, data-driven approach to deriving finer-grained and more comprehensive aspects.
@@ 3.1 Definition of aspect
While aspects are outlined in review guidelines and used in related work, a formal definition of aspect is lacking. To address this, we propose an operational definition of aspect as a characteristic of a paper that a reviewer makes a judgment on when evaluating the paper, which is later used to compare the manuscripts to each other and to the quality standards provided by the review guidelines. From this definition, we do not consider terms such as Acceptance Decision as aspects, since they are not characteristics of a paper. An aspect can be general, such as Soundness, or specific, such as Missing Citations on Controlled Generation. However, aspects that are either too general or too specific are difficult to use for comparing manuscripts to each other or to the publication standards. Therefore, an aspect taxonomy that accommodates different levels of granularity is needed.
@@ 3.2 Method
To establish a comprehensive set of aspects, we selected paper reviews from NLP and machine learning (ML) conferences across different time periods. We randomly selected 50 papers from each of the NLPeer Dycke et al. (2023) and EMNLP23222Publicly available through the OpenReview API. datasets. We used the keywords “natural” and “language” to filter NLP-related papers from ICLR. We randomly selected 50 NLP-related papers from each of the ICLR conferences from 2020 to 2024. We selected 350 papers in total, corresponding to 1094 reviews. We segmented the reviews into sentences using NLTK Punkt Tokenizer Bird and Loper (2004) and performed the identification at the sentence level. See Appendix B for more details.
@@
We used OpenAI GPT-4o [^3] GPT-4o-2024-08-06, from Oct 31 to Nov 20, 2024. to identify aspects from the reviews. The prompt we used is shown in Table 7. Since the identification was performed in an unsupervised setting, the identified aspects appear inconsistent, with variations like Result and Results. We post-processed the results to reduce such variations. We first identified the most frequently occurring aspects in the results, which were used as keywords to categorize and match the remaining ones. We grouped related terms together, such as Clarification and Clarity. We removed terms that we considered to be too general, such as Weakness, Strength, Question, and Comment. For the terms that cannot be matched, we omitted those that appear less than 50 times in the annotations.
We notice that our post-processing method may lead to unrelated terms being grouped together. For example, using Improvement as a keyword groups Performance Improvement and Improvement Recommendation (i.e., suggestions to improve paper quality) together. We manually checked the results to verify and correct inappropriate groupings.
In cases where a single term contains multiple aspects, it is placed into more than one category (e.g., Comparison with Related Work is included in both Comparison and Related Work categories).
@@ 3.3 Results
Some aspects are not present in our taxonomy as review forms for major NLP venues have dedicated fields for their evaluation, making them appear much less frequently in the review text (e.g., Reproducibility). See Table 8 for examples of the dataset we created which pairs aspects with real-life reviews and Table 9 for the aspect frequencies.
@@ 3.4 Validity check
We conducted a human evaluation to verify the model annotations, involving three human annotators to evaluate the LLM annotations mapped to the coarse label set. We follow Yuan et al. (2022), asking the annotators to determine whether a review sentence addresses the aspects identified by GPT-4o (see Table 14 for examples). We observe that on average human annotators agree with 91% of the LLM annotations, along with fair inter-annotator agreement as measured by Fleiss’ Kappa Fleiss (1971) (see Table 3). These results suggest that the model annotations largely align with human judgments, and that human annotators can effectively understand our taxonomy.
See Appendix B.2 for more details.
@@ 4.1 Method
It is important to note that PAP is inherently more challenging than RAP. PAP operates as a heuristic method grounded in statistics, or in a more advanced setting (i.e., when the input is the full paper), goes beyond heuristics to infer strengths or weaknesses in certain aspects of the paper. In contrast, RAP resembles summarization, as it extracts aspects directly from the review text, making it comparatively less challenging than PAP.
We modeled both tasks as multi-label sequence classification. For both tasks, we tested bag-of-words with random forest (BoW+RF) and RoBERTa Liu et al. (2019). Both models are strong in supervised settings, and they are lighter-weight alternatives to LLMs. We used focal loss Lin et al. (2017) to address label imbalance. For PAP, we also experimented with GPT-4o in both zero-shot and few-shot settings. Our evaluation metrics include precision, recall, F1 score, and Jaccard similarity score. See Appendix C for more details.
@@ 5.1 Aspect analysis
Table 5: The tracks with the 5 highest frequencies of Analysis in EMNLP23. Table 22 shows the full results.
@@ 5.2 Review comparison
To compare the reviews, we predicted the aspects in each review and calculated the Jaccard similarity between sets of aspects. Figure 2 visualizes the similarity between each pair of human-written and LLM-generated reviews.
@@
Figure 2: The heatmap of the Jaccard similarity between each pair of the human-written reviews and LLM-generated reviews generated using EMNLP23 papers and Liang et al. (2024b)’s prompt. Figure 6 in Appendix D.2 shows the rest of the results.
We obverve that LLM-generated reviews show a higher degree of similarity to each other in terms of aspects. This suggests that LLM-generated reviews may be more generic than human-written ones, with the model tending to comment on similar sets of aspects across different papers. This could also imply that the quality of LLM-generated reviews is still lacking in terms of specificity.
@@ 5.3 LLM-generated review detection
Table 6 shows the detection results. We used Jaccard similarity and compared our method with an implementation of Equation 2 using Sentence-BERT (SBERT) Reimers and Gurevych (2019), which calculates the similarity between the embeddings of review texts. Experiments were conducted using both the coarse and fine label sets, as well as ACL’18 which is a commonly used aspect set in previous studies (see Table 1). We calculated a random baseline that selects a review at random as the LLM-generated review. Note that the fine aspect set consistently outperform ACL’18 and SBERT, and our method is robust across LLM-generated reviews generated using different prompts.
@@
While this detection approach does not achieve the same level of performance as the reference-based and zero-shot approaches, such as the 90% accuracy reported by Yu et al. (2024) or our own GPT-4o results in Table 24, it nevertheless provides valuable insights. It demonstrates that current LLMs tend to generate reviews that are generic in terms of aspects. A key advantage of this approach lies in interpretability, opening new opportunities for broader applications such as review quality assessment. Moreover, the performance gains observed when using the fine label set (see Table 6) highlight the importance of aspect granularity–finer labels reveal patterns that are useful in distinguishing LLM-generated reviews from human-written ones, patterns that coarse labels fail to capture. Thus, we consider this approach promising, and leave the development of better performing aspect-based detectors to future work.
5.4 Recommendations
Based on our results, we make the following recommendations on selecting aspect granularity. The coarse label set is more suitable for high-level analysis tasks, such as analyzing review focus across different tracks and venues (Section 5.1). The fine label set is more appropriate for tasks that require nuanced analysis, where capturing specific and detailed feedback is critical, such as review comparison (Section 5.2) and LLM-generated review detection (Section 5.3). In some cases, a hybrid strategy might be the best option. For example, when evaluating review quality, one can first apply the coarse label set to assess coverage (i.e., whether key aspects are addressed), and then use the fine label set to assess specificity of the review. Aspect granularity is a design choice, and the optimal configuration depends on the task and user needs.
@@ 6 Conclusion
In this paper, we have introduced a data-driven approach to peer review aspect analysis. We have provided an operational definition of aspect, and developed an semi-automatic approach to identify aspects from peer reviews. We have proposed a taxonomy that involves a comprehensive set of aspects with different granularity, and introduce a new dataset of peer reviews augmented with aspects. We introduced two tasks, paper aspect prediction and review aspect prediction, and have shown how they contribute to a detailed empirical study of aspects. Our results demonstrate that fine-grained, data-driven aspects complement coarse aspects from review guidelines, and allow for more nuanced review comparison and new interpretable approaches for LLM-generated review detection.
@@
As discussed in Section 3.1, aspect is difficult to define. We adopted an operational definition as a practical approach. Defining individual aspects is also challenging. For instance, it is difficult to determine the scope of “Methodology.” Polysemy further complicates this issue–for example, “Improvement” may refer to either a method’s improvement, or places where the paper can be improved. We did a manual inspection of the categorization results to minimize the impact of these issues. While our study provides a foundation for data-driven research on review aspect, future work may seek to refine and expand upon our current definition.
The taxonomy was constructed by an expert annotator (one of the authors of this paper) with expertise in NLP and peer review, based on domain knowledge. This taxonomy may not be optimal, and alternative approaches to categorizing aspects may exist. Our taxonomy is an attempt to streamline the analysis and application of aspects, and it has been shown to be effective for the purposes of this study. We deem a future multiple-expert study in fine-grained aspect identification promising.
@@
Consistency is a fundamental issue when working with LLM-generated annotations. As reported in Section 3.4, GPT-4o indeed shows some sensitivity to prompts and temperature. Our validity checks suggest a reasonable degree of consistency and reliability of the LLM annotations. The applications of these aspect annotations, especially in the LLM-generated review detection task, provide further support for their reliability. Though we specify the seed (See Appendix B and C), the exact reproducibility of the results related to the closed-weights LLMs like GPT-4o is a concern. As open models become more capable, experimentation with alternative open models and aspect schemata will become possible.
@@
For human annotation, we recognize that using aspect labels from our taxonomy may introduce potential label bias. A straightforward solution to this would be to ask the annotators to do an open-ended annotation task without pre-defined labels. However, such study is very challenging in terms of the cognitive burden on annotators and ensuring annotation consistency. Automation bias is also a concern. While we address this by asking annotators to provide explanations for their annotations, follow-up work can explore alternative experimental strategies to measure and mitigate automation bias, e.g., through the use of distractor items.
@@
In addition, we point at potential selection bias due to data availability. All papers associated with the ICLR20-24 and EMNLP23 datasets are camera-ready versions. This may affect the validity of the results of PAP experiments in Section 4 since in practice PAP deals with (potentially lower-quality) submission manuscripts. We do not consider the use of camera-ready versions problematic for review comparison and LLM-generated review detection in Section 5.2 and 5.3. A further bias in item selection might be introduced by data imbalance with respect to paper acceptance. The EMNLP23 dataset contains only 9 rejected papers; in NLPeer, 69% of the papers are accepted papers Dycke et al. (2023), which does not correspond to a natural distribution of submissions. Such skew may lead to an overestimation of aspect frequency: aspects commonly associated with accepted papers may appear more frequently than they would in a more balanced dataset. Consequently, some of our findings may be influenced by this bias. If these overrepresented aspects are used to inform review forms or guidelines, they may introduce new biases into the review process, which reduces their effectiveness. As more review data become available, this limitation can be mitigated.
@@
Finally, in Section 5.3, we only focus on the commonly used end-to-end general prompts for our review generator. We did not consider prompts that involve paper-specific aspects, as we consider these to be human-in-the-loop prompts, where the user must first read the paper to identify paper-specific aspects and then incorporate them into the prompt, which would require an experimental setup beyond our scope. This approach represents a form of human-AI collaboration, and we plan to explore this direction in future work.
@@
This work does not suggest or imply that the proposed task of predicting the aspects to focus on a given paper may replace human involvement. Instead, it is designed to assist reviewers by recommending aspects to consider. Reviewers retain full autonomy over their decisions regarding whether to include the suggested aspects. Since all the review data used in this study are publicly available and anonymized, processing it with commercial LLMs such as OpenAI GPT-4o does not raise ethical concerns.